Write an equation of the line (3,7) and parallel to y=3x−8. Write the equation in the form x=​a, y=​b, or y=mx+b.
Answer by ankor@dixie-net.com(22740)

Write an equation of the line (3,7) and parallel to y=3x−8.
Write the equation in the form x=​a, y=​b, or y=mx+b.
:
Parallel lines have the same slope to right away we know the equation is:
y = 3x + b
We know line has a point x=3, y=7, so we can write it
7 = 3(3) + b
7 = 9 + b
now we can find b
7 - 9 = b
-2 = b
therefore the equation is
y = 3x - 2